OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? My answer is 240 Am I right?

OpenStudy (anonymous):

Im getting 230 but I could of messed a step up..

OpenStudy (anonymous):

how did you get your answer?

OpenStudy (anonymous):

230 isn't one of my options its either A: 576 B:240 C:144 Or D:960

OpenStudy (anonymous):

Lol im really confused on this!

OpenStudy (anonymous):

what work did you do to get the answer?

OpenStudy (anonymous):

Oh wait, I just got 576

OpenStudy (anonymous):

1x = 0.6x + 384 1x - 0.6x = 384 0.4x = 384 0.4x/0.4 = 384/0.4 x = 960 960 x 0.6 = 576

OpenStudy (anonymous):

Does that look right?